OpenStudy (anonymous):

Medal and fan to first person to answer 3/5(15-8)-1/10 a.11/10 b.21/10 c.31/10 d.41/10

OpenStudy (anonymous):

d 41\10

OpenStudy (anonymous):

can you help on couple more questions

OpenStudy (anonymous):

shure!

OpenStudy (anonymous):

2(1/2+3/4)+6 a.57/8 b.19/2 c.17/2 d.22/3

OpenStudy (anonymous):

c. 17/2

OpenStudy (anonymous):

(12-4) + 8/2 a.4 b.6 c.8 d.10

OpenStudy (anonymous):

that would be 12? but it's not on there 12-4=8 8/2=4 8+4=12

OpenStudy (anonymous):

ok so what would it be then

OpenStudy (anonymous):

12, it comes out to be 12, and I sugjest using this as proof,and for further math questions http://www.wolframalpha.com/input/?i=%2812-4%29+%2B+8%2F2

OpenStudy (anonymous):

I checked it this thing is the best calculator I know of.

OpenStudy (anonymous):

ok thanks for ur help

OpenStudy (anonymous):

any time :D